OpenStudy (anonymous):

What is the distance between (2, 0) and (5, 8)?

OpenStudy (anonymous):

\[d=\sqrt{(x _{2}-x _{1})^2+(y _{2}-y _{1})^2}\] plug the values into the formula

OpenStudy (anonymous):

= sqrt ( 9+36) =sqrt(9*5) = 3sqrt(5)